A. For each statement below, select T if the statement is True, F if it is false and NA if there is not enough information to determine if it is true or false.

Based on the spectrum, it is possible to determine the presence of a nitrogen atom.        T / F / NA

Based on the spectrum, it is possible to determine the presence of an oxygen atom.       T / F / NA

Based on the spectrum, it is possible to determine the presence of a sulfur atom.            T / F / NA

Based on the spectrum, it is possible to determine the presence of a chlorine atom.        T / F / NA

Based on the spectrum, it is possible to determine the presence of a bromine atom.        T / F / NA

B. Which atom or group has been fragmented off in the base peak?

C. Why is there a peak at 88? What does this peak correspond to?

D. What is the structure of the compound?

MASS SPECTRUM 49 (100%) (99) 10 60 Rel. Abundance %E1) 98 40 51 (31%) 47 (13.9%) 20.- (%) 88 0 0 20 40 60 80 100 m/z
Answer #1

(A)  

NA

As per nitrogen rule if even number of nitrogen present the molecular ion will be of even mass number, off number of nitrogen causes molecular ion of odd mass number.  

(D)

We find peaks at 84,86,88

Are due to presence of two chlorine atoms.

For both 35 Cl m/z = 84

For both 37 Cl m/z = 88

For one 35 Cl and one 37 Cl , m/z = 86

Compound is dicloromethane

CH2Cl2,